### v6

@@ 4.1 Preliminary lemmas
This confirms that $y$ is a mild solution of Eq. 1.1, and it is evident that Eq. 3.1 holds. The uniqueness of the mild solution can be proved similarly as in [17, Theorem 1.1].
@@
Finally, combining Eqs. 2.10, 4.23 and 4.24 establishes the regularity assertions in Eqs. 3.4 and 3.5. Moreover, the regularity assertion in Eq. 3.2 follows directly from Eq. 3.4 and the observation that $G\in C([0,T];L^{q})$. This completes the proof of Theorem 3.1.
@@
We first summarize some standard results as follows.
@@
Let $p,q\in(1,\infty)$. For any $g_{h}\in L^{p}(0,T;X_{h})$, the following inequality holds for all $\epsilon>0$:
$\max_{t\in[0,T]}\|S_{0,h}g_{h}(t)\|_{\dot{H}_{h}^{2-\frac{2}{p}-\epsilon,q}}\leqslant c\|g_{h}\|_{L^{p}(0,T;L^{q})}.$
@@
For any $t\in[0,T]$, recognizing that $S_{0,h}g_{h}(t)=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s$, we proceed as follows:
$\|S_{0,h}g_{h}(t)\|_{\dot{H}_{h}^{2-\frac{2}{p}-\epsilon,q}}$ $=\left\|\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s\right\|_{\dot{H}_{h}^{2-2/p-\epsilon,q}}$
$\leqslant\int_{0}^{t}\left\|e^{(t-s)\Delta_{h}}g_{h}(s)\right\|_{\dot{H}_{h}^{2-2/p-\epsilon,q}}\,\mathrm{d}s$
@@
where step (i) invokes Eq. 4.5 with $\alpha=2-2/p-\epsilon$, and step (ii) uses Hölder’s inequality. The uniform boundedness of the integral
@@
with respect to $t\in[0,T]$ then establishes the lemma. ∎
@@
Assume that $v\in L^{\infty}$. Let $y$ and $y_{h}$ be the mild solutions of Eqs. 1.1 and 2.7, respectively, and let $G$ and $G_{h}$ be defined by Eqs. 2.11 and 2.12, respectively. For simplicity, let $z:=y-G$ and $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z$. Then, for any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\varepsilon>0$, the following inequalities hold:
@@
$\mathbb{E}\int_{0}^{T}\lVert{(y-P_{h}y)(t)}\rVert_{L^{4}}^{p}\lVert{y(t)}\rVert_{L^{8}}^{2p}\,\mathrm{d}t\leqslant ch^{2-\varepsilon}.$ (4.29)
@@
$\mathbb{E}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\leqslant ch^{2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p},$ (4.30)
$\mathbb{E}\int_{0}^{T}\lVert{\xi_{h}(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\leqslant ch^{p+2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p-\varepsilon}.$ (4.31)
@@
Now let us prove Eq. 4.30. By invoking Eq. 4.2 and replicating the reasoning behind [18, Equation (2.2)], we establish that
$\mathbb{E}\int_{0}^{T}\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{\dot{H}^{1,2}}^{2}\,\mathrm{d}t$
@@
$\mathbb{E}\int_{0}^{T}\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{L^{6}}^{2}\,\mathrm{d}t.$
@@
$\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{\frac{4\beta}{3\beta-3}}\leq\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{L^{6}}^{2},$
@@
$\mathbb{E}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t$
@@
Invoking the inequality Eq. 4.25 with $p$ replaced by $4p\beta(4\beta-(3\beta-3)p)$, we find
$\mathbb{E}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\leqslant ch^{2-\frac{3}{2}(1-\frac{1}{\beta})p}.$
Given that Theorem 3.1 ensures $y$ resides within $L^{\frac{4\beta}{3\beta-3}}(\Omega\times(0,T);L^{2\beta})$, we can mirror the preceding argument to deduce that
$\mathbb{E}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\leqslant ch^{2-\frac{3}{2}(1-1/\beta)p}.$
The inequality Eq. 4.30 is thus substantiated by combining the above derived inequalities.
Finally, by invoking Eq. 4.27 and adhering to the same logical progression as demonstrated in the proof of Eq. 4.30, we deduce Eq. 4.31. This concludes the proof. ∎
@@
where $G$ and $G_{h}$ are defined by Eqs. 2.11 and 2.12, respectively. From Eq. 1.1 and the fact $G^{\prime}(t)=\Delta G(t)$ for $t\in[0,T]$, we deduce that
@@
$\mathbb{E}\lVert{\eta_{h}(t)}\rVert_{L^{2}}^{p}$ $=p\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Big{[}\Delta_{h}\eta_{h}(s)+\xi_{h}(s)+\eta_{h}(s)$
$\qquad\qquad\qquad\qquad\qquad{}+(G_{h}-P_{h}G)(s)+\big{(}y^{3}-y_{h}^{3}\big{)}(s)\Big{]}\,\mathrm{d}\mu\,\mathrm{d}s$
$\qquad{}+\frac{p}{2}\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s$
$\qquad{}+\frac{p(p-2)}{2}\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\lambda_{k}\Big{[}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}\mu\Big{]}^{2}\,\mathrm{d}s$
@@
$I_{1}$ $:=p\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Delta_{h}\eta_{h}(s)\,\mathrm{d}\mu\,\mathrm{d}s,$
$I_{2}$ $:=p\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot(\xi_{h}+\eta_{h}+G_{h}-P_{h}G)(s)\,\mathrm{d}\mu\,\mathrm{d}s,$
$I_{3}$ $:=p\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}y^{3}-(P_{h}y)^{3}\big{]}(s)\,\mathrm{d}\mu\,\mathrm{d}s,$
$I_{4}$ $:=p\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}(P_{h}y)^{3}-y_{h}^{3}\big{]}(s)\,\mathrm{d}\mu\,\mathrm{d}s,$
$I_{5}$ $:=\frac{p}{2}\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s,$
$I_{6}$ $:=\frac{p(p-2)}{2}\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\lambda_{k}\Big{[}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}\mu\Big{]}^{2}\,\mathrm{d}s.$

### v7

@@ 4.1 Preliminary Lemmas
This confirms that $y$ is a mild solution of Eq. 1.1, and it is evident that Eq. 3.1 holds. To prove the uniqueness of the mild solution, assume $\widetilde{y}$ is another mild solution. Consider the difference $e=y-\widetilde{y}$. Then $e$ satisfies the equation
$\begin{cases}\mathrm{d}e(t)=(\Delta e+e+\widetilde{y}^{3}-y^{3})(t)\,\mathrm{d}t+\big{(}F(y(t))-F(\widetilde{y}(t))\big{)}\,\mathrm{d}W_{H}(t),&t\in[0,T],\\ e(0)=0.\end{cases}$
By an argument similar to that used in Eq. 4.14, we find that
$\lVert{e}\rVert_{L^{2}(\Omega\times(0,T);L^{2})}\leqslant 0,$
which implies that $e=0$ almost surely, confirming the uniqueness of the mild solution to Eq. 1.1.
@@
Finally, combining Eqs. 2.9, 4.16 and 4.17 establishes the regularity assertions in Eqs. 3.4 and 3.5. Moreover, the regularity assertion in Eq. 3.2 follows directly from Eq. 3.4 and the observation that $G\in C([0,T];L^{q})$. This completes the proof of Theorem 3.1.
@@
We begin by summarizing several well-established results. As known from [8], it holds that
$\|P_{h}\|_{\mathcal{L}(L^{q})}\text{ is uniformly bounded with respect to }h\text{ for all }q\in(1,\infty).$ (4.18)
Consequently, the operator $P_{h}$ possesses the following approximation property:
$\|I-P_{h}\|_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant ch^{\alpha}\quad\text{for all }\alpha\in[0,2]\text{ and }q\in(1,\infty).$ (4.19)
Inequality (2.5) from [1], combined with the observation that
$\lVert{\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{1,2})}\leqslant 1,$
yields, through interpolation, that for any $q\in[2,\infty)$, the operator norm
$\lVert{\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{1,q})}$
is uniformly bounded with respect to $h$. Utilizing this fundamental stability estimate, the approximation property Eq. 4.19, the interpolation theory as stated in [22, Theorem 2.6], and the fact that $\Delta^{-1}$ is a bounded linear operator mapping from $\dot{H}^{-s,q}$ to $\dot{H}^{2-s,q}$ for all $s\in[0,2]$, standard arguments then lead to the conclusion that
$\|P_{h}\Delta^{-1}-\Delta_{h}^{-1}P_{h}\|_{\mathcal{L}(\dot{H}^{\beta-2,q},\dot{H}^{\alpha,q})}\leqslant ch^{\beta-\alpha}\quad\text{for all }0\leqslant\alpha\leqslant 1\leqslant\beta\leqslant 2.$ (4.20)
Let $e^{t\Delta_{h}}$, for $t\in[0,\infty)$, denote the analytic semigroup generated by the discrete Laplace operator $\Delta_{h}$. Using the established estimates for the resolvents of $\Delta_{h}$ from [1] and applying (2.128) from [38], we obtain the well-known growth estimate: for any $\alpha\in[0,1]$, $q\in(1,\infty)$, and $t\in(0,T]$,
$\lVert{(-\Delta_{h})^{\alpha}e^{t\Delta_{h}}}\rVert_{\mathcal{L}(L^{q})}\leqslant ct^{-\alpha}.$ (4.21)
Furthermore, utilizing techniques analogous to those found in Theorem 6.10 and Lemma 20.7 of [35], one can readily derive the following well-known inequality: for any $\alpha\in[0,2]$, $q\in(1,\infty)$, and $t\in[0,T]$,
$\lVert{e^{t\Delta_{h}}P_{h}-P_{h}e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}+\lVert{(I-P_{h})e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant c\min\Big{\{}h^{\alpha},\frac{h^{2}}{t^{1-\alpha/2}}\Big{\}}.$ (4.22)
For the special case where $q=2$, the reader is referred to Theorem 3.2 of [35] for a detailed treatment of this result.
Let $q\in(1,\infty)$. For any $\alpha\in\mathbb{R}$, denote by $\dot{H}_{h}^{\alpha,q}$ the space $X_{h}$ equipped with the norm
$\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}:=\|(-\Delta_{h})^{\alpha/2}v_{h}\|_{L^{q}}\quad\text{for all }v_{h}\in X_{h}.$
According to Theorem 3.1 (ii) in [17], the boundedness of the purely imaginary powers of $-\Delta_{h}$ in $\mathcal{L}(\dot{H}_{h}^{0,q})$ is independent of the spatial mesh size $h$. Consequently, by applying Theorem 4.17 from [22], we deduce that for any $\alpha\in(0,2)$,
$c_{0}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\leqslant\|v_{h}\|_{[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}}\leqslant c_{1}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\quad\text{for all }v_{h}\in X_{h},$ (4.23)
where $c_{0}$ and $c_{1}$ are positive constants independent of $h$. Here, $[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}$ denotes the complex interpolation space between $\dot{H}_{h}^{0,q}$ and $\dot{H}_{h}^{2,q}$ (see [22, Chapter 2]). This result is important as it enables the application of interpolation theory to the space $\dot{H}_{h}^{s,q}$ for all $s\in\mathbb{R}$. For example, given the standard inverse estimate
$\|v_{h}\|_{\dot{H}_{h}^{2,q}}\leqslant ch^{-2}\|v_{h}\|_{L^{q}}\quad\text{for all }v_{h}\in X_{h},$
we can invoke [22, Theorem 2.6] to derive the following inverse estimate: for any $0\leqslant\alpha<\beta<\infty$,
$\|v_{h}\|_{\dot{H}_{h}^{\beta,q}}\leqslant ch^{\alpha-\beta}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\quad\text{for all }v_{h}\in X_{h}.$
Furthermore, for any $\alpha\in[0,1]$, there exist positive constants $c_{0}$ and $c_{1}$, independent of $h$, such that
$c_{0}\|v_{h}\|_{\dot{H}^{\alpha,q}}\leqslant\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\leqslant c_{1}\|v_{h}\|_{\dot{H}^{\alpha,q}}\quad\text{for all }v_{h}\in X_{h}.$
Henceforth, the aforementioned theoretical results will be utilized without further explicit reference.
We now introduce a discrete deterministic maximal $L^{p}$-regularity estimate and its stochastic counterpart. For a function $g_{h}\in L^{1}(0,T;\dot{H}_{h}^{0,2})$, the discrete deterministic convolution operator $S_{0,h}$ is defined as
$(S_{0,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s,\quad\text{for }t\in[0,T].$ (4.24)
For a process $g_{h}\in L_{\mathbb{F}}^{2}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,2}))$, the discrete stochastic convolution operator $S_{1,h}$ is given by
$(S_{1,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}W_{H}(s),\quad\text{for }t\in[0,T].$ (4.25)
The proofs of the following two lemmas can be found in [16] for the first and in [17] for the second.
@@
Suppose $p,q\in(1,\infty)$. For any $g_{h}\in L^{p}(0,T;X_{h})$, the following discrete deterministic maximal $L^{p}$-regularity estimate is valid:
$\Big{\lVert}{\frac{\mathrm{d}}{\mathrm{d}t}S_{0,h}g_{h}}\Big{\rVert}_{L^{p}(0,T;\dot{H}_{h}^{0,q})}+\lVert{S_{0,h}g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{2,q})}\leqslant c\lVert{g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{0,q})}.$
Lemma 4.8.
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. For any $g_{h}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))$, the following discrete stochastic maximal $L^{p}$-regularity estimate holds:
[…]
Additionally, for the discrete deterministic convolution operator $S_{0,h}$, we establish the following property.
Lemma 4.9.
Let $p,q\in(1,\infty)$. For any $g_{h}\in L^{p}(0,T;\dot{H}_{h}^{0,q})$, the following inequality holds for all $\epsilon>0$:
$\|S_{0,h}g_{h}(t)\|_{C([0,T];\dot{H}_{h}^{2-2/p-\epsilon,q})}\leqslant c\|g_{h}\|_{L^{p}(0,T;L^{q})}.$
@@
For any $t\in[0,T]$, from the identity $S_{0,h}g_{h}(t)=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s$, we deduce that
$\|S_{0,h}g_{h}(t)\|_{\dot{H}_{h}^{2-2/p-\epsilon,q}}$ $\leqslant\int_{0}^{t}\left\|e^{(t-s)\Delta_{h}}g_{h}(s)\right\|_{\dot{H}_{h}^{2-2/p-\epsilon,q}}\,\mathrm{d}s$
@@
where step (i) uses the estimate Eq. 4.21 with $\alpha=2-2/p-\epsilon$, and step (ii) applies Hölder’s inequality. Since the integral
@@
is uniformly bounded with respect to $t\in[0,T]$, the desired inequality follows, thereby completing the proof of the lemma. ∎
We proceed to present some auxiliary estimates.
@@
Suppose that $v\in L^{\infty}$. Consider the mild solution $y$ to equation (1.1), the strong solution $y_{h}$ to equation (2.11), and the function $G$ defined by equation (2.10). Additionally, let $G_{h}(t):=e^{t\Delta_{h}}P_{h}v$ for $t\in[0,T]$. For notational simplicity, let $z:=y-G$ and $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z$. For any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\varepsilon>0$, the following inequalities hold:
@@
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(y-P_{h}y)(t)}\rVert_{L^{4}}^{p}\lVert{y(t)}\rVert_{L^{8}}^{2p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\varepsilon}.$ (4.31)
@@
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p},$ (4.32)
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{\xi_{h}(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\bigg{]}\leqslant ch^{p+2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p-\varepsilon}.$ (4.33)
Here, $\beta^{\prime}$ denotes the conjugate exponent of $\beta$.
@@
Now let us prove Eq. 4.32. By invoking Eq. 4.18 and replicating the reasoning behind [19, Equation (2.2)], we establish that
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{\dot{H}^{1,2}}^{2}\,\mathrm{d}t\bigg{]}$
@@
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{L^{6}}^{2}\,\mathrm{d}t\bigg{]}.$
@@
$\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{\frac{4\beta}{3\beta-3}}\leqslant\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{L^{6}}^{2},$
@@
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\bigg{]}$
@@
Invoking the inequality Eq. 4.27 with $p$ replaced by $4p\beta/(4\beta-(3\beta-3)p)$, we find
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3}{2}(1-\frac{1}{\beta})p}.$
Given that Theorem 3.1 ensures $y$ belongs to $L^{\frac{4\beta}{3\beta-3}}(\Omega\times(0,T);L^{2\beta})$, we can replicate the previous argument to conclude that
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3}{2}(1-1/\beta)p}.$
Hence, Eq. 4.32 follows by combining the aforementioned inequalities. To deduce Eq. 4.33, we rely on Eq. 4.29 and proceed analogously to the proof of Eq. 4.32. This completes the proof. ∎
@@
where $G$ is defined by Eq. 2.10, and $G_{h}$ is given by
$G_{h}(t):=e^{t\Delta_{h}}P_{h}v,\quad t\in[0,T].$ (4.36)
From Eq. 1.1 and the fact $G^{\prime}(t)=\Delta G(t)$ for $t\in[0,T]$, we deduce that
@@
$\mathbb{E}\lVert{\eta_{h}(t)}\rVert_{L^{2}}^{p}$ $=p\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Big{(}\Delta_{h}\eta_{h}(s)+\xi_{h}(s)+\eta_{h}(s)$
$\qquad\qquad\qquad\qquad\qquad{}+(G_{h}-P_{h}G)(s)+\big{(}y^{3}-y_{h}^{3}\big{)}(s)\Big{)}\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]}$
$\qquad{}+\frac{p}{2}\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s\biggr{]}$
$\qquad{}+\frac{p(p-2)}{2}\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\lambda_{k}\Big{(}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}\mu\Big{)}^{2}\,\mathrm{d}s\biggr{]}$
@@
$\allowdisplaybreaks I_{1}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Delta_{h}\eta_{h}(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
$I_{2}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot(\xi_{h}+\eta_{h}+G_{h}-P_{h}G)(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
$I_{3}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}y^{3}-(P_{h}y)^{3}\big{]}(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
$I_{4}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}(P_{h}y)^{3}-y_{h}^{3}\big{]}(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
$I_{5}$ $:=\frac{p}{2}\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s\biggr{]},$
$I_{6}$ $:=\frac{p(p-2)}{2}\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\lambda_{k}\Big{(}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}\mu\Big{)}^{2}\,\mathrm{d}s\biggr{]}.$